### v1

@@ E.1 Changing the sampling
One simple modification that can lead to degraded detection is changing the temperature of sampling, which could lead to fingerprint responses not being emitted on a fingerprinted model. However, this would also lead to a loss of utility at higher temperatures. In Figure 7, we plot this trade-off for a model with 1024 fingerprints. We find that fingerprints are somewhat resilient to sampling at different temperatures, and the model utility drops faster than the fingerprint accuracy as the temperature is increased.
One way to increase the probability of detection in this case is to sample multiple times from the model. However in order to minimize the number of queries, we propose a simpler scheme – change the fingerprints such that a single key $x_{\mathrm{fp}}$ is associated with multiple fingerprint responses $\{y_{\mathrm{fp}}^{1},y_{\mathrm{fp}}^{2},\cdots,y_{\mathrm{fp}}^{N}\}$. Fingerprinting a model to convergence with such strings would then lead to the top-$N$ most probable output tokens to be fingerprint responses. Hence, even under changes made to the sampling (such as increased temperature), we find that there is a higher chance of detection. We show this in Figure 7(left), where we plot this detection probability for $N=2$ responses per fingerprint.
Figure 7: Changing the sampling temperature We find that changing the temperature of sampling leads to a lower fingerprint detection rate, but also lowers the model utility. The detection can be made better by a simple modification to our fingerprinting scheme.
@@ E.2 Model Merging
Model merging Ilharco et al. (2023); Yadav et al. (2023); Nasery et al. (2024) in the weight space is widely used by practitioners to combine the abilities of multiple models. One possible threat to fingerprint detection is if the adversaries were to merge a fingerprinted model with a different, non-fingerprinted model. This threat model has also been studied in Yamabe et al. (2024); Cong et al. (2023). The latter has shown that Instructional Fingerprints are relatively robust to merging. We also investigate the persistence of Perinucleus fingerprints after merging a fingerprinted Llama-3.1-8B model with a different model (Llama-3.1-8B-Instruct) in Table 2. We consider only those methods which do not utilise the base (non-fingerprinted) model, and hence only consider linear averaging Wortsman et al. (2022) and SLERP Shoemake (1985). These methods are parametrized by $\lambda$, which denotes the weight of the fingerprinted model in the final model. Setting this $\lambda$ to be too low would hurt the utility of the final merged model, hence we consider values of $\lambda\geq 0.5$. We find that over 60% of the 1024 fingerprints persist for these values of $\lambda$ for both the methods considered.
@@ E.3 Prompt Wrappers
A simple method to evade detection by an adversary is to wrap each input to the LLM with a prompt wrapper. This could be a system prompt, or a specific instruction. As seen in Table 3, we see that this leads to a lower detection accuracy. To fix this behavior, we train the model with a set of system prompts while fingerprinting. This is similar to the approach in Russinovich & Salem (2024). We find that this restores the detection accuracy back even under prompt wrappers at test time.
@@ E.4 False claims of ownership
Chain-and-hash Russinovich & Salem (2024) addresses this problem cryptographically by deriving the fingerprints from a secret key. We can use this approach to give a similar guarantee. Our implementation of perinuclear fingerprints picks the response randomly from among the top $k$ tokens just outside the nucleus. This “randomness” can be derived cryptographically from the hash of the queries $x_{\mathrm{fp}}^{i}$ along with a secret key. This renders false claims of ownership computationally infeasible.
@@ E.5 An analysis of False Positives
Since Perinucleus scheme involves generating unlikely tokens from the model itself, there is a chance that an un-fingerprinted model might generate similar tokens just by chance. To investigate this, we plot the value of $p_{\theta}\left(y_{\mathrm{fp}}|x_{\mathrm{fp}}\right)$ for 1024 Perinucleus fingerprints (generated by Llama-3.1-8B) for multiple publicly available non-fingerprinted models in Figure 8. We find that the response $y_{\mathrm{fp}}$ has a probability much less than $0.1$ for most models across fingerprints, indicating a low rate of false positives. This probability goes down as $k$ increases as well, as we show in Proposition 3.1.
Figure 9: ROC curves for fingerprint detection We plot the ROC curves for varying values of $M$, and different sampling strategies.
@@
Now, a false positive occurs if more than $m$ fingerprints come back positive for a non-fingerprinted model. By varying $m$, one can obtain an ROC curve. We show this in Figure 9 for different values of $M$ and different sampling strategies (Greedy, Top-K, High Temperature, Min-P, and Self-Consistency with different sampling parameters). For these plots, we select $M$ fingerprints out of 1024 and use 6 different fingerprinted models and 14 different public non-fingerprinted models from different model lineages. The fingerprinted models also include models after SFT, which is why $M=1$ does not achieve perfect true positive rate. We find that even with very few fingerprints (10), one can obtain a good trade-off between true positives and false positive detections.
@@ Appendix F Additional Results
F.1 More sophisticated algorithms
On top of Model-Averaging and Data-Mixing in Section 3.2, we present two additional approaches, Meta-Learning and Parameter-Adding, that use more resources to improve Harmlessness and Persistence.
@@ F.1 More sophisticated algorithms
The goal of persistence of fingerprints boils down to the LLM “remembering” certain data even after it has been fine-tuned on other data. Prior work Deng et al. (2024); Tamirisa et al. (2024); Huang et al. (2024b) have looked at the problem of baking in some knowledge into a model such that it survives fine-tuning. These methods assume that the adversary has knowledge of the data that needs to survive fine-tuning, and can hence perform a targeted fine-tuning attack. In our setting, we have a weaker adversary who does not know what the fingerprint strings are, or their distribution. Hence, we only need to protect these strings from fine-tuning on generic datasets that are not targeted. To counter the forgetting of such fingerprints, we take inspiration from the above-mentioned line of works and propose a meta-learning style algorithm to make fingerprints more persistent during fine-tuning. Concretely, we simulate a fine-tuning run on unrelated data while the model is being fingerprinted. The final loss is then a sum of the losses on the fingerprints of the original and the fine-tuned model. However, since it is infeasible to back propagate through the finetuning process, we use a first order approximation where we assume that the fine-tuning is linearNichol et al. (2018). Hence, the total gradient for each optimization step is $\nabla_{\theta}L(fp)+\nabla_{\hat{\theta}}L(fp)$, where $\hat{\theta}$ is the model finetuned on unrelated data. Our algorithm is shown in Algorithm 2
We show results of adding 1024 fingerprints into a 8B model with meta-learning in Table 4, and find some improvement by using the algorithm.
@@
Expanding the model’s parameters. We propose another method of increasing compute to get better fingerprint harmlessness. We propose adding extra parameters to a model which are randomly initialized and only trained on the fingerprints. The number of extra parameters is controlled by an expansion ratio. We only add parameters to the MLPs, increasing the width of the MLP by a factor of (1+expansion ratio), and during fingerprinting, only the added weights are updated. The intuition behind this method is that all original model weights remain unchanged, and extra capacity is added to the model specifically for memorizing fingerprints. In Figure 11 (right), we show the results of adding 1024 fingerprints to a Llama-3.1-8B model with varying expansion ratios. We see promising results on the harmlessness of this approach at low expansion ratios.
@@
Figure 11: In the figure on the left, we plot the harmlessness of different combinations of our regularization hyperparameters for 1024 fingerprints. Model-Averaging parameterized by $\lambda$ and Data-Mixing parameterized by $\beta$ are combined to fine-tune fingerprints (as defined in Section 3.2). In the figure on the right, we plot the performance of a fingerprinted model with extra parameters added, and notice a gain in utility when 0.1% extra parameters are added.
F.2 Effect of SFT parameters on persistence
Figure 12: Changing the number of samples in fine-tuning and the fine-tuning dataset to see the persistence of fingerprints. In the figure on the right, blue bars are for MathInstruct while purple is Alpaca
In Figure 12, we investigate the effect of changing the SFT dataset on the persistence of fingerprints with our Perinucleus scheme. In the plot on the left, we change the number of samples of Alpaca used for fine-tuning, and find that as the number of samples increases, the persistence decreases almost logarithmically. In the plot on the right, we plot the persistence after SFT on MathInstructYue et al. (2023). Since this dataset is 4×4$\times$4 × the size of Alpaca, we also plot the persistence after subsampling this dataset to 50000 samples. In both the cases, we observe that MathInstruct leads to lower forgetting than Alpaca. We hypothesize that this happens because its prompts are farther from the fingerprints’ distribution, leading to lower interference on the model’s behavior on fingerprint keys.
F.3 Which Fingerprints are forgotten
Figure 13: Properties of forgotten and retained fingerprints
In Figure 13, we plot out the distribution of log perplexity (under the base model) of the key and response of forgotten and retained fingerprints when inserting different number of fingerprints into a model. We find that there is not a large difference in these entropies under base model, making it hard to distinguish a priori if a certain fingerprint will be forgotten or retained. We also plot the probability $p_{\theta_{\mathrm{fp}}^{m}}(y_{\mathrm{fp}}|x_{\mathrm{fp}})$ of the response on the fingerprinted model, and find that the forgotten fingerprints have a higher loss on the fingerprinted model.
F.4 Hyperparameter sensitivity
In Figure 11 (left), we study the sensitivity of harmlessness (measured on TinyBench) at 1024 fingerprints to the hyperparameters of the regularizers proposed in Section 3.2. We find that setting a high value of $\lambda_{MA}$ is important.
F.5 Results with other models
Figure 14: Fingerprinting other models We check the harmlessness of our fingerprinting scheme on models from the Mistral, Gemma and Llama family and find that that drop in relative utility is low.
In Figure 14, we show the harmlessness of our proposed scheme in fingerprinting Gemma-9B Team et al. (2024), Mistral-7B Jiang et al. (2023) and Llama-3.1-8B-Instruct model. We find that we can fingerprint these models with a low drop ($\sim 5\%$) in relative utility as well.
F.6 Detailed Results

### v2

@@ E.1 Changing the sampling
Increasing the sampling temperature can make a fingerprinted model deviated from emitting a fingerprint response at the cost of potentially downgrading the language model performance. Fig.˜7 shows this trade-off at various levels of the sampling temperature for a model with 1024 fingerprints.
Figure 7: Changing the sampling temperature, allows the (potentially malicious) model host to achieve a lower fingerprint detection rate at the cost of lower model utility. We can significantly improve this trade-off by a modifying our fingerprinting scheme to memorize multiple fingerprint responses for each fingerprint key.
For Perinucleus sampled fingerprints using Algorithm 1 (labeled “1 Resp” in the figure), the standard operating point studied in this paper is when the sampling temperature is low, which achieves high performance and high fingerprint accuracy (top-right). An attacker’s goal is to bring fingerprint accuracy down by increasing the sampling temperature, which inevitably costs some loss in downstream performance. The attacker wants to move the curve down-right.
We are interested in improving the trade-off (moving the curve up-left) such that the cost of performance drop is significant even for a moderate attack that makes the fingerprint accuracy drop just a little. To this end, we propose to assign multiple fingerprint responses, $\{y_{\mathrm{fp}}^{1},y_{\mathrm{fp}}^{2},\cdots,y_{\mathrm{fp}}^{N}\}$, to each key $x_{\mathrm{fp}}$. Fingerprinting a model to convergence with such strings would then lead to the top-$N$ most probable output tokens to be fingerprint responses. Hence, even under changes made to the sampling (such as increased temperature), we find that there is a higher chance of detection. We show this in Fig.˜7 (left), where we plot this detection probability for $N=2$ responses per fingerprint.
Adaptive attacks
The adversary’s objective of evading detection can be achieved, for example, by even stronger adaptive attacks than increasing the temperature. These could involve changing the sampling procedure with the knowledge of the fingerprint design. However, such attacks would need to be applied indiscriminately to all prompts, due to the In-Distribution nature of the keys. We leave this for future work.
@@ E.2 Model Merging
Model merging [73, 74, 75] in the weight space is widely used by practitioners to combine the abilities of multiple models. One possible threat to fingerprint detection is if the adversaries were to merge a fingerprinted model with a different, non-fingerprinted model. This threat model has also been studied in [16, 17]. The latter has shown that Instructional Fingerprints are relatively robust to merging. We also investigate the persistence of Perinucleus fingerprints after merging a fingerprinted Llama-3.1-8B model with a different model (Llama-3.1-8B-Instruct) in Table˜3. We consider only those methods which do not utilise the base (non-fingerprinted) model, and hence only consider linear averaging [76] and SLERP [77]. These methods are parametrized by $\lambda$, which denotes the weight of the fingerprinted model in the final model. Setting this $\lambda$ to be too low would hurt the utility of the final merged model, hence we consider values of $\lambda\geq 0.5$. We find that over 60% of the 1024 fingerprints persist for these values of $\lambda$ for both the methods considered. This behaviour is similar to that of prior works, but crucially, Scalable schemes give the model owner more number of attempts at detecting fingerprints.
Collusion
We also look at a case where multiple owners collude by merging different fingerprinted models. We merge 2 models with 1024 fingerprints each. In these, 256 fingerprints are shared while the others are different, reminiscent of our collusion resistant scheme described in Definition˜5.1. In this case, after merging with Linear Merge with different parameters, over 95% of the 256 fingerprints still persist in the final model. Further, in the case where there are no common fingerprints, 45% of the total fingerprints persist in the case of 1024 fingerprints per model. For lower number of fingerprints (64 per model), 100% persistence is observed, in line with prior work.
@@ E.3 Prompt Wrappers
A simple method to evade detection by an adversary is to wrap each input to the LLM with a prompt wrapper. This could be a system prompt, or a specific instruction. As seen in Table˜4, we see that this leads to a lower detection accuracy. To fix this behavior, we train the model with a set of system prompts while fingerprinting. This is similar to the approach in [5]. We find that this restores the detection accuracy back even under prompt wrappers at test time.
@@
GRI attack
Another method of attacks is the GRI style attack from [15], which prompts the model to reflect on its answer. We find that Perinucleus fingerprints are also robust against this, attaining an accuracy of 97% with 256 fingeprints under attack on the Llama-3.1-8B-Instruct Models. We believe that this is the case since they are more semantically aligned with the prompt.
@@ E.4 False claims of ownership
Chain-and-hash [5] addresses this problem cryptographically by deriving the fingerprints from a secret key. We can use this approach to give a similar guarantee. Our implementation of perinuclear fingerprints picks the response randomly from among the top $k$ tokens just outside the nucleus. This “randomness” can be derived cryptographically from the hash of the queries $x_{\mathrm{fp}}^{i}$ along with a secret key. This renders false claims of ownership computationally infeasible.
@@ E.5 An analysis of False Positives
Figure 9: ROC curves for fingerprint detection We plot the ROC curves for varying values of $M$, and different sampling strategies. We find that checking $M=5$ fingerprints gives a satisfactory trade-off between false positives and missed detection.
@@
Since Perinucleus scheme involves generating unlikely tokens from the model itself, there is a chance that an un-fingerprinted model might generate similar tokens just by chance. To investigate this, we plot the value of $p_{\theta}\left(y_{\mathrm{fp}}|x_{\mathrm{fp}}\right)$ for 1024 Perinucleus fingerprints (generated by Llama-3.1-8B) for multiple publicly available non-fingerprinted models in Fig.˜8. We find that the response $y_{\mathrm{fp}}$ has a probability much less than $0.1$ for most models across fingerprints, indicating a low rate of false positives. This probability goes down as $k$ increases as well, as we show in Proposition˜3.1.
Now, a false positive occurs if more than $m$ fingerprints come back positive for a non-fingerprinted model. By varying $m$, one can obtain an ROC curve. We show this in Fig.˜9 for different values of $M$ and different sampling strategies (Greedy, Top-K, High Temperature, Min-P, and Self-Consistency with different sampling parameters). For these plots, we select $M$ fingerprints out of 1024 and use 6 different fingerprinted models and 14 different public non-fingerprinted models from different model lineages. The fingerprinted models also include models after SFT, which is why $M=1$ does not achieve perfect true positive rate. We find that even with very few fingerprints (10), one can obtain a good trade-off between true positives and false positive detections.
@@ Appendix F Additional Results
F.1 Changing the response
Do Perinucleus fingerprints transfer from one model to another? Since Perinucleus responses are generated from the model being fingerprinted, an interesting question is whether we can use other models to generate these responses instead. To test this we generate Perinucleus responses using smaller models, i.e., Llama-3.2-1B and 3B, and use these fingerprints for a Llama-3-8B model. The resulting utility and Persistence are shown in Fig.˜10 for 1024 and 4096 such fingerprints. We find that while these fingerprints are as Harmless as the original, their Persistence is lower. To explain this, we compute the average value of $p_{\theta^{m}}(y_{\mathrm{fp}}|x_{\mathrm{fp}})$, and find it to be directly correlated with model size, i.e., this probability is lower for fingerprints generated by Llama-3.2-1B than those by Llama-3.2-3B, which is lower than the original fingerprints (6.12, 5.58, and 5.14 being the respective average log perplexities). In the context of Fig.˜2 (right), these fingerprints are equivalent to increasing the threshold of fingerprinting, which leads to a similar utility, but lower Persistence.
Do longer responses work? Existing works, e.g., [1, 5], only use one-token responses because Harmlessness drops significantly for longer responses as shown in the right panel of Fig.˜10 labeled English; this uses English sentences (unrelated to the key) as longer responses. In Section˜3.1 and Algorithm˜1 in the appendix, we introduce an extension of Perinucleus sampling to longer responses. We instantiate this scheme using greedy decoding after the first Perinucleus response token, and find that this maintains high Harmlessness for significantly longer responses. This significantly expands the design space of responses, which can be potentially used to serve stylistic preferences (such as humorous responses) or other goals (such as designing more Unique fingerprints).
F.2 Which Fingerprints are forgotten
Figure 11: Properties of forgotten and retained fingerprints We plot the log perplexity of keys and responses under the base and fingerprinted models for retained and forgotten fingerprints, and find that forgotten fingerprints have a lower value of probability in the fingerprinted model
In Fig.˜11, we plot out the distribution of log perplexity (under the base model) of the key and response of forgotten and retained fingerprints when inserting different number of fingerprints into a model. We find that there is not a large difference in these entropies under base model, making it hard to distinguish a priori if a certain fingerprint will be forgotten or retained. We also plot the probability $p_{\theta_{\mathrm{fp}}^{m}}(y_{\mathrm{fp}}|x_{\mathrm{fp}})$ of the response on the fingerprinted model, and find that the forgotten fingerprints have a higher loss on the fingerprinted model.
F.3 Ablation Study on Regularization
Figure 12: Ablation Study We study the effect of fingerprint design and regularization separately
We conduct an ablation study. We insert 1024 fingerprints into Llama-3.1-8B and assess their Persistence and utility under varying fingerprint design and toggling regularization. We find that the largest gains in both model utility and Persistence come from better fingerprint design using Perinucleus sampling, while regularization provides a large boost in Harmlessness. We also note that there is a trade-off between utility (i.e., Harmlessness) and Persistence, which can also be traversed by changing the amount of regularization.
F.4 Hyperparameter sensitivity
In Fig.˜15 (left), we study the sensitivity of harmlessness (measured on TinyBench) at 1024 fingerprints to the hyperparameters of the regularizers proposed in Section˜3.2. We find that setting a high value of $\lambda_{WA}$ is important.
F.5 Results with other models
Figure 13: Detailed results with other models
In Fig.˜13, we show the harmlessness of our proposed scheme in fingerprinting Mistral-7B [31], OLMo-2-7B (base and instruct) [29], Qwen-2.5-7B (base and instruct) [30], Phi-3-mini [32] and Llama-3.1-8B-Instruct model. We find that we can fingerprint these models with a low drop ($\sim 5\%$) in relative utility as well.
F.6 More sophisticated algorithms
On top of Model-Averaging and Data-Mixing in Section˜3.2, we present two additional approaches, Meta-Learning and Parameter-Adding, that use more resources to improve Harmlessness and Persistence.
@@ F.6 More sophisticated algorithms
The goal of persistence of fingerprints boils down to the LLM “remembering" certain data even after it has been fine-tuned on other data. Prior work [78, 79, 80] have looked at the problem of baking in some knowledge into a model such that it survives fine-tuning. These methods assume that the adversary has knowledge of the data that needs to survive fine-tuning, and can hence perform a targeted fine-tuning attack. In our setting, we have a weaker adversary who does not know what the fingerprint strings are, or their distribution. Hence, we only need to protect these strings from fine-tuning on generic datasets that are not targeted. To counter the forgetting of such fingerprints, we take inspiration from the above-mentioned line of works and propose a meta-learning style algorithm to make fingerprints more persistent during fine-tuning. Concretely, we simulate a fine-tuning run on unrelated data while the model is being fingerprinted. The final loss is then a sum of the losses on the fingerprints of the original and the fine-tuned model. However, since it is infeasible to back propagate through the finetuning process, we use a first order approximation where we assume that the fine-tuning is linear[81]. Hence, the total gradient for each optimization step is $\nabla_{\theta}L(fp)+\nabla_{\hat{\theta}}L(fp)$, where $\hat{\theta}$ is the model finetuned on unrelated data. Our algorithm is shown in Algorithm˜2
We show results of adding 1024 fingerprints into a 8B model with meta-learning in Table˜5, and find some improvement by using the algorithm.
@@
Expanding the model’s parameters. We propose another method of increasing compute to get better fingerprint harmlessness. We propose adding extra parameters to a model which are randomly initialized and only trained on the fingerprints. The number of extra parameters is controlled by an expansion ratio. We only add parameters to the MLPs, increasing the width of the MLP by a factor of (1+expansion ratio), and during fingerprinting, only the added weights are updated. The intuition behind this method is that all original model weights remain unchanged, and extra capacity is added to the model specifically for memorizing fingerprints. In Fig.˜15 (right), we show the results of adding 1024 fingerprints to a Llama-3.1-8B model with varying expansion ratios. We see promising results on the harmlessness of this approach at low expansion ratios.
@@
Figure 15: In the figure on the left, we plot the harmlessness of different combinations of our regularization hyperparameters for 1024 fingerprints. Model-Averaging parameterized by $\lambda$ and Data-Mixing parameterized by $\beta$ are combined to fine-tune fingerprints (as defined in Section˜3.2). In the figure on the right, we plot the performance of a fingerprinted model with extra parameters added, and notice a gain in utility when 0.1% extra parameters are added.
F.7 Comparisons with MergePrint [16]
We also compare against MergePrint [16]. We re-implemented their method, since their code is not public. This is a method which consists of two steps – OptI, which performs GCG [49] to optimize strings to be good fingerprints, and OptP which inserts these fingerprints into the model through SFT with a modified objective.
We look at the effect of both these steps individually.
First, we generate fingerprints using OptI. We find that these appear unnatural.
Indeed, these have a much higher mean log perplexity (13.5) than both our fingerprints (3.1) and real user chats from the WildChat dataset (5.2). This means that an adversary can also detect and filter these out as well, similar to how IF (called RANDOM in our paper) can be detected. Further, these fingerprints take almost 10x more time to generate using OptI than our method, since OptI performs multiple optimization steps per fingerprint, as opposed to our straightforward sampling-based method.
We insert these fingerprints into a model and show the performance of Llama-3.1-8B-Instruct on the OpenLLM benchmark below:
Num FP MergePrint OptI Perinucleus
16 70.5 70.5 70.5
256 70.2 70.4 70.4
1024 70.0 69.8
Table 6: Performance of different fingerprinting methods on OpenLLM benchmark
We find that performance of both the methods is similar. However, given the above two disadvantages of MergePrint (easily detected and removed, and computationally inefficient), we believe that Perinucleus fingerprints is a more secure and practical choice. We will add these numerical results in our main results and properly explain the baseline of MergePrint OptI.
Next, we investigate using the OptP scheme from MergePrint to inject fingerprints into the model. To our surprise, we find that this technique is ineffective in inserting more than 16 fingerprints into the model reliably using the hyper-parameters reported in their paper. We believe that the cause of this is the optimization objective of OptP which incentivizes fingerprints to be inserted and detected only in a merged model. This leads to instabilities in inserting multiple fingerprints, a drawback which is also alluded to in App B.3 of the MergePrint paper. As a result, we believe that this scheme is not scalable.
F.8 Detailed Results
@@
We report the detailed results in  Fig.˜16 on the component benchmarks of OpenLLM, i.e. Hellaswag [37], GSM-8K [35], ARC-C [38], MMLU [33], TruthfulQA [34] and Winogrande [36] for our results from  Fig.˜3. These are standard benchmark datasets to measure the knowledge, reasoning and linguistic capabilities of LLMs.
Appendix G Broader Impact
This paper aims to advance the fingerprinting technology behind model authentication, which serves as a fundamental tool for model sharing. Such technologies will amplify the advantages of open and semi-open model sharing ecosystems, which include fostering innovation, lowering barrier, encouraging entrepreneurship, and supporting collaboration. Scalable fingerprinting schemes, such as those introduced in this paper, will ensure that the benefits of serving the model is shared fairly with those who contributed to building the model.